OpenStudy (mikezack123):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = 2/3x − 10 y = 2/3x + 10 y = −3/2x − 10 y = −3/2x + 10

OpenStudy (anonymous):

If its perpendicular, then the slope is is the negative reciprocal of the original slope

OpenStudy (anonymous):

so if the slope was 5x then the perpendicualr slope is -1/5x

OpenStudy (mikezack123):

so the answer is A?

OpenStudy (anonymous):

nope first you gotta get it into y=mx+b form

OpenStudy (anonymous):

can you solve for y?

OpenStudy (mikezack123):

yes Y=mx+b

OpenStudy (mikezack123):

i mean B is the answer

OpenStudy (anonymous):

i meant in the above equation its not in the form Ax+By=C

OpenStudy (mikezack123):

B is totes the answer right?

OpenStudy (anonymous):

not quite you gotta do a little work first here: 2x − 3y = 3

OpenStudy (mikezack123):

how about C

OpenStudy (anonymous):

-3y=-2x+3 y=2/3x-1

OpenStudy (anonymous):

now m=slope has to be the opposite reciprocal, do that

OpenStudy (anonymous):

and then pllug the slop in, with the point, in the point slope formula which is y-y1=m(x-x1) where y1=2 and x1=-8 and when you solve for y, you get your answer

OpenStudy (mikezack123):

ca you solve for y for me idk how 2 do it

OpenStudy (anonymous):

I'm here to help not answer but what is the negative reciprocal of 2/3?

OpenStudy (mikezack123):

whats reciprocal

OpenStudy (anonymous):

the reciprocal of \[\frac{ a }{ b } is =\frac{b}{a}\]

OpenStudy (mikezack123):

so the answer is -3/2

OpenStudy (anonymous):

yes now with m=-3/2 y1 being the y coordinat of (-8,2) and x1 being the x coordiante of it, plug the slope and the point into the point slope formula and tell me what you get (the other x and y dont change) \[y-y1=m(x-x1)\]

OpenStudy (mikezack123):

the answer is C

OpenStudy (anonymous):

you got it

OpenStudy (mikezack123):

woo!